OpenStudy (anonymous):

find the equation of the tangent line to the circle (x+4)^2+(y+1)^2=25 at (-1,3)

OpenStudy (anonymous):

differentiate this equation wrt 'x' you'll get { 2(x+4)+2(y+4)dy/dx=0} put x= -1 and y=3 then calculate dy/dx .. now you have one point and slope of line . can determine the eq. If not use (y-3)=dy/dx (x+1) . you'll get eq . of tangent line.

OpenStudy (anonymous):

so whats the answer ?

OpenStudy (anonymous):

3x+7y=18